OpenStudy (anonymous):

Find the balance in the account. $700 principal earning 5%, compounded monthly, after 9 years (A)$79,380.00 (B)$1,085.93 (C)$16,088.19 (D)$1,096.79

OpenStudy (tkhunny):

i = 0.05 -- Annual Interest Rate j = I/12 -- Monthly Interest Rate r = 1+j -- Monthly accumulation factor Begin: 700 Month1: 700r Month2: 700r^2 Month3: 700r^3 ... Month 72: ??